OpenStudy (anonymous):

Using complete sentences, explain the two ways to simplify sqrt(180) and provide the simplified form.

OpenStudy (tkhunny):

Oh, where, oh where is my prime factorization? I findeth it beneath the radical. What joy to see factors of 2. Pull them out until you are through. 180 2*90 2*2*45 Alas, 2 is done, perhaps 3 will give some? 2*2*45 2*2*3*15 2*2*3*3*5 No more threes, but what have we left? Only a 5 for the radical to heft. \(\sqrt{2*2*3*3*5} = \sqrt{2\cdot 2}\cdot\sqrt{3\cdot 3}\cdot \sqrt{5}\) "Is there more," we ask? Certainly, a simplify will complete the task. \(\sqrt{2\cdot 2}\cdot\sqrt{3\cdot 3}\cdot \sqrt{5} = 2\cdot 3\cdot\sqrt{5}\) Yet, only one thing we lack, then delight we return to the question's author. \(2\cdot 3\cdot\sqrt{5} = 6\sqrt{5}\) It just says "complete sentences". It doesn't say you have to be boring and technical. Branch out and have a little fun with it!